# On the complexity of covering points by disjoint segments and by guillotine cuts

### v1

@@ Abstract
We show that two geometric cover problems in the plane, related to covering points with disjoint line segments, are NP-complete. Given $n$ points in the plane and a value $k$, the first problem asks if all points can be covered by $k$ disjoint line segments; the second problem treats the analogous question for $k$ guillotine cuts.
@@ 1 Introduction
In this note, we are interested in two variants of the problem of covering points with lines. In the classic point line cover problem, one is given $n$ points in the plane, with coordinates $(x_{1},y_{1}),\dots,(x_{n},y_{n})$, and the goal is to find a set of lines $\ell_{1},\dots,\ell_{r}$ of minimum cardinality such that for each $1\leq i\leq n$, point $(x_{i},y_{i})$ lies on at least one line $\ell_{j}$, for some $1\leq j\leq r$. As mentioned above, the point line cover problem has been known to be NP-hard for decades [6].
Our interest is on variants of the point line cover problem where the objects used to cover points are disjoint. Perhaps the most natural variant in this direction is covering with disjoint line segments, which will be our main focus. In addition, we also consider the problem of covering points with guillotine cuts. Informally, a guillotine cut is a straight line segment (possibly degenerating into a halfline or line) that extends from one segment to another segment. Guillotine cuts (also known as guillotine partitions) are subdivisions formed by recursively applying guillotine cuts. They appear naturally in cutting stock problems, and the resulting segments are disjoint, except at endpoints. See Figure 1 for an illustration of both problems.
@@
Surprisingly, to the best of our knowledge, these natural variants of the point cover problem have not been studied before. Less surprisingly, we prove in this paper that these problems are NP-complete. Proving this is the contribution of this short note. While not very surprising, confirming NP-completeness is an important prerequisite for any further algorithmic study of these natural problems.
@@ 2 Covering points by disjoint segments
In this section present an NP-completeness proof for the problem of covering points with disjoint line segments.
@@ 4 Conclusions
We have shown that two variants of the problem of covering points by lines, where there is an additional disjointedness requirement, result in NP-complete problems. In particular, we considered the problems of covering points by disjoint line segments, and covering points by guillotine cuts (which are technically not disjoint, but they contain no proper segment intersections). Having established the—rather unsurprising—complexity of the problem justifies exploring other interesting avenues. In particular, future work will focus on understanding whether algorithms for non-disjoint versions of the point line cover problem can be applied to the disjoint version, and on proving bounds between solution sizes of the different variants.

### v2

# On the complexity of covering points by guillotine cuts

@@ Abstract
We show that the problem of covering a set of points in the plane with a minimum number of guillotine cuts is NP-complete. To that end, first we present a new NP-completeness proof for the problem of covering points with disjoint line segments. Then, we adapt the proof to show that the problem remains NP-complete when the segments are guillotine cuts.
@@ 1 Introduction
In this note, we focus on two variants of the problem of covering points with lines. In the classic point line cover problem, one is given $n$ points in the plane, with coordinates $(x_{1},y_{1}),\dots,(x_{n},y_{n})$, and the goal is to find a set of lines $\ell_{1},\dots,\ell_{r}$ of minimum cardinality such that for each $1\leq i\leq n$, point $(x_{i},y_{i})$ lies on at least one line $\ell_{j}$, for some $1\leq j\leq r$. As mentioned above, the point line cover problem has been known to be NP-hard for decades [7].
Our interest is on variants of the point line cover problem where the objects used to cover points are disjoint. Perhaps the most natural variant in this direction is covering with disjoint line segments. Surprisingly, the complexity of this problem remained open until recently, when Grelier [4] proved that the problem is NP-hard.
The main goal of this work is to prove a similar result for the problem of covering points with guillotine cuts. Informally, a guillotine cut is a straight line segment (possibly degenerating into a halfline or line) that extends from one segment to another segment. Guillotine cuts (also known as guillotine partitions) are subdivisions formed by recursively applying guillotine cuts. They appear naturally in cutting stock problems, and the resulting segments are disjoint, except at endpoints. See Figure 1 for an illustration of both problems.
While the problem of covering points by disjoint segments is known NP-hard, the reduction used by Grelier [4] is from the problem Maximum Independent Set in Intersection Graphs of Segments. Since it is not evident how to adapt such a reduction for the particular case of segments that are guillotine cuts, in this work we present a new proof of NP-hardness for covering points by disjoint segments, reducing from planar monotone 3-SAT, which we can later adapt to work with guillotine cuts.
While verifying that these problems are NP-hard is not very surprising, confirming hardness is an important prerequisite for any further algorithmic study of the problems.
@@ 2 Covering points by disjoint segments
In this section present a new NP-completeness proof for the problem of covering points with disjoint line segments, using a construction that we can later adapt to guillotine cuts.
@@ 4 Conclusions
We have shown that the problem of covering points by guillotine cuts is NP-complete. In particular, first we revisited the problem of covering points by disjoint line segments, and presented a new hardness proof based on planar monotone 3-SAT. A new proof seemed necessary as an intermediate step, given that the adaptation of the previous proof from disjoint line segments [4] to guillotine cuts did not look straightforward. Having established the—rather unsurprising—complexity of the problem justifies exploring other more interesting avenues. In particular, future work will focus on understanding whether existing algorithms for variants of the line cover problem can be applied to the guillotine cut version, and on proving bounds between solution sizes of the different variants.